OpenStudy (anonymous):

aarong please help me with my chemistry homework find the volume of a 93.91g sample of air, which has a density of 0.00130g/ml

OpenStudy (abb0t):

density is defined as \(\sf \color{}{\frac{mass}{volume}}\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

0.00130g/ml equals 98.91 over volume!! but I got .122083 is that right

OpenStudy (abb0t):

I don't have a calculator, but as long as you divided, you should be correct.

OpenStudy (anonymous):

so 0.00130g/ml divided by 98.91 will get you the answer..

OpenStudy (anonymous):

i mean 93.91 g

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

but my teachers answer sheets says 72200 mL

OpenStudy (anonymous):

and i got a different answers help

OpenStudy (abb0t):

You probably put it in the calcuator wrong, because it should be that answer. Recheck

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

my teacher or me place get a calculator so you can help me.. Don't understand my teacher

OpenStudy (abb0t):

You are plugging it incorrectly!

OpenStudy (anonymous):

0.00130 divided by 93.91 equal.. Thats what i did or do i touch 2nd 0.00130 divided by 93.91 equals..

OpenStudy (abb0t):

93.91 divided by 0.00130!!

OpenStudy (anonymous):

i see what i did wrong u told me to divided. you didn't tell me what to divded or how to/ its mass divided by the density..

OpenStudy (anonymous):

i got the answer d equals 42.7 over 21.13 whats the density?

OpenStudy (abb0t):

You always start with what you're given, so u can cancel out the units, like from Algebra: \(\sf \color{}{a \times \frac{b}{a}}\) the a's cancel outyou're left with "b"

OpenStudy (anonymous):

d equals 42.7 over 21.13 whats the density?

OpenStudy (anonymous):

Help

OpenStudy (aaronq):

what is your question?

OpenStudy (anonymous):

Will sample float in water that has a mass of 42.7g and a volume of 21.13mL?

OpenStudy (anonymous):

I got arronq helping me so bye abb0t i really didn't understandy what you were teaching me sorry..

OpenStudy (anonymous):

your like my teacher, your not great a teaching okay i'm just saying that's why i found this website. So i can be helped by anyone okay. I'm sorry

OpenStudy (anonymous):

now arronq is gone and he was going to help me with my question...

OpenStudy (anonymous):

SOMEONE HELP ME PLEASE